OpenStudy (anonymous):

if a+b is not equal to 0 and a-b is not equal to 0 then (a^2-b^2)/(a^2+2ab+b^2) - (ab-a^2)/(a^2-b^2) is equal to what?

OpenStudy (anonymous):

factorize it first.

OpenStudy (anonymous):

agreed...^^^

OpenStudy (anonymous):

um.... i got ab(a-b)/ ab(a+2+b) - a(b-a)/ ab(a-b)

OpenStudy (anonymous):

\(\large \frac{a^2-b^2}{a^2+2ab+b^2 } - \frac{ab-a^2}{a^2-b^2}\) = \(\large \frac{\cancel{(a+b)}(a-b)}{\cancel{(a+b)}(a+b) } - \frac{a\cancel{(b-a)}^{-1}}{(a+b)\cancel{(a-b)}}\) = \(\large \frac{a-b}{a+b}+\frac{a}{a+b} \) = \(\large \frac{2a-b}{a+b} \)

OpenStudy (anonymous):

oh thanks :)

OpenStudy (anonymous):

yw...:)